OpenStudy (anonymous):

A vase contains 12 tulips and 20 roses. What is the ratio of roses to the total number of flowers in the vase? A:3 to 8 B:3 to 5 C:5 to 8 D:5 to 3

OpenStudy (texaschic101):

there are 20 roses there are (20 + 12) = 32 total flowers roses : flowers = 20 to 32 or 20/32 which reduces to 5/8 or 5 to 8

OpenStudy (anonymous):

So it is 5:8?Thank yall so much!yall are life saviors!!!

OpenStudy (anonymous):

lol thanx yall!

OpenStudy (texaschic101):

happy to help. Just remember, ratios can be written as fractions as well. Ex. 6 to 10 = 6/10.....2 to 3 or 2:3 = 2/3